### v2

# Hyper-Compression: Model Compression via Hyperfunction

Hyper-Compression: Model Compression via Hyperfunction
Feng-Lei Fan E-mail: zeng@math.cuhk.edu.hk Juntong Fan E-mail: zeng@math.cuhk.edu.hk Dayang Wang Feng-Lei Fan, Juntong Fan, and Dayang Wang contribute equally. E-mail: zeng@math.cuhk.edu.hk Jingbo Zhang E-mail: zeng@math.cuhk.edu.hk Zelin Dong E-mail: zeng@math.cuhk.edu.hk Shijun Zhang E-mail: zeng@math.cuhk.edu.hk Ge Wang E-mail: zeng@math.cuhk.edu.hk Tieyong Zeng E-mail: zeng@math.cuhk.edu.hk
(December 14, 2024)
[…]
The rapid growth of large models’ size has far outpaced that of GPU memory. To bridge this gap, inspired by the parsimonious relationship between genotype and phenotype, we turn the model compression problem into the issue of parameter representation to propose the so-called hyper-compression. The hyper-compression uses a hyperfunction to represent the parameters of the target network per ergodic theory, that addresses the following approximation problem: if a low-dimensional dynamic system can fill the high-dimensional space eventually. Empirically, the proposed hyper-compression enjoys the following merits: 1) Preferable compression ratio; 2) No post-hoc retraining; 3) Affordable inference time; and 4) Short compression time. It compresses LLaMA2-7B in an hour and achieves close-to-int4-quantization performance, without retraining and with a performance drop of less than 1%. Our work can facilitate the harmony between the scaling law and the stagnation of hardware upgradation in terms of saving both computation and data. We have open-sourced our code for readers’ free download and evaluation.
1 Introduction
Recently, due to the pursuit of the scaling law, the escalating demand for memory and computational resources by large models has presented formidable challenges for their deployment in resource-constrained environments [8], particularly with the stagnation of hardware Moore’s Law [25]. Besides technical obstacles, the substantial cost associated with deployment also causes the dominance of major institutions in the community, which introduces risks of integrity, bias, security, and other critical aspects. Thus, how to serve large models well is not only a technological imperative but also a strategic pursuit for social good.
One prevalent way to serve large models is to develop effective model compression approaches that crop the size of large models while maintaining the performance of the original model. Currently, the landscape of model compression predominantly revolves around four different classes of algorithms: pruning, quantization, knowledge distillation, and low-rank decomposition [34] whose fundamental frameworks were established years ago. However, the compression efficacy of these methods is challenging to scale. For instance, even the most extreme case of quantization, represented by 1-bit integers, remains unexciting because the compression ratio is at best a constant, which is handicapped to address the enlarged gap between models and the hardware resource. Moreover, when compressing a model a lot, the model’s output is severely distorted. Thus, retraining and recalibration are mandated to recover the model’s performance.
Here, we introduce hyper-compression, a novel and general-purpose approach that redefines model compression as a problem of parameter representation. This concept stems from hyper-networks that are smaller networks but can generate weights for a significantly larger target network, in analogy to the biological observation that a comparatively small genome can encode a large neuronal network. We generalize the concept of hypernets into what we term a ‘hyperfunction’ (HF) and hypothesize that weights of a large network can be encoded into a small parametric function. Mathematically, we use a parametric function $W_{n}=h(n;\Theta)$ to encode the relationship between locations and weights of the target network, where $W_{n}$ is the $n$-th parameter of the target network, and $\Theta$ is the hyperparameters of $h$. Should the memory footprint of $\Theta$ be significantly less than that of $W$, the function $h$ can be stored on devices, allowing for the dynamic hierarchical querying of weights during inference and thus facilitating the efficient deployment of large-scale models.
One of the most important problems that ergodic theory investigates is if a low-dimensional dynamic system can eventually visit a high-dimensional space given sufficient time. We empirically find that the ergodic theory indicates a suitable hyperfunction and leads to a performant hyper-compression algorithm that offers several distinct advantages, succinctly summarized as PNAS: 1) Preferable compression ratio; 2) No post-hoc retraining&recalibration; 3) Affordable inference time; and 4) Short compression time. More favorably, our method is compatible with other compression methods such as pruning to further escalate the compression efficacy multiplicatively. Lastly, we underscore that model compression via ergodic theory is a novel view whose intrinsic mechanism, relating to chaotic theory, Kolmogorov theorem, Diophantine representation problem, etc., remains unclear. Given the criticality of model compression, we report our empirical results first, and exploring the mechanism of hyper-compression will be our future work.
Result and Analysis
This section illustrates the efficacy of our novel compression methodology on three widely utilized large, middle, and small models: LlaMA2-7B [24], UNet [20], and MobileNetV3 [15]. Distinctively, our compression technique does not necessitate post-hoc retraining, even at high compression ratios. For example, we achieve up to a 7.87×7.87$\times$7.87 × reduction in the model size of UNet within $1\%$ performance drop. This is a significant improvement over traditional methods that often require extensive retraining to restore efficacy. This capability sets a new benchmark in model compression. Detailed descriptions of our algorithm and experiments are provided in Supplementary Information.
Table 1: The compression effectiveness of our method on UNet, MobileNetV3, and LlaMA2-7B. As a reference, the compression ratios of INT4 quantization in LlaMA-7B, UNet, and MobileNetV3 are 4×4$\times$4 ×, 8×8$\times$8 ×, and 8×8$\times$8 ×, respectively.
Model File Size (GB) Average (%)
LlaMA2-7B [24] 12.50 65.88
Sheared-LlaMA-1.3B [28] 5.01 (2.50$\times$) 50.31
TinyLlaMA [31] 4.09 (3.06$\times$) 51.53
LiteLlaMA [^1] For more details, please visit the Hugging Face website of LiteLlaMA: https://huggingface.co/ahxt/LiteLlama-460M-1T. 0.86 (14.53$\times$) 40.41
LlaMA-7B + HF 4.80 (2.60$\times$) 64.89
Sheared-LlaMA-1.3B + HF 0.98 (12.76$\times$) 49.76
TinyLlaMA + HF 0.78 (16.03$\times$) 50.65
LiteLlaMA + HF 0.39 (32.05$\times$) 39.72
Model File Size (MB) Dice (%)
UNet [20] 51.10 99.86
Pruning 20.30 (2.53$\times$) 96.34
HF 6.49 (7.87$\times$) 99.71
Pruning + HF 2.88 (17.74$\times$) 96.45
Model File Size (MB) Top-1 Accuracy (%)
MobileNetV3 [15] 5.95 74.41
Pruning 2.22 (2.68$\times$) 69.32
HF 1.65 (3.61$\times$) 73.93
Pruning + HF 0.47 (12.66$\times$) 68.47
Table 2: Timing results for different batch sizes on two devices using two methods with a UNet network.
Batch Size 4060 A40
Original (s) Ours (s) Original (s) Ours (s)
1 4.35 4.43 1.50 2.06
2 3.95 4.49 1.44 2.05
4 3.63 4.77 1.45 2.22
8 3.67 7.94 1.44 2.46
Table 3: The compression time of our method on UNet, MobileNetV3
and LlaMA2-7B
Model Compression Time (s)
UNet + HF [20] 30.00
MobileNetV3 + HF 11.63
Model Compression Time (min)
LlaMA2-7B + HF 53
Sheared-LlaMA-1.3B + HF 7
Preferable compression ratio. As shown in Table 4, we apply the same pruning method used in Sheared-LlaMA-1.3B, TinyLlaMA, and LiteLlaMA, combined with our compression technique. We evaluate our compressed models on eight downstream tasks: 0-shot accuracy on SciQ, WinoGrande, ARC-E, 25-shot ARC-C, 10-shot HellaSwag, 32-shot BoolQ, NQ, and 5-shot MMLU. The average test results of eight downstream tasks are shown in Table 4, where “File Size" indicates the size of the file that stores the model. The comparison of individual tasks is put in Supplementary Information. Notably, our method can compress LlaMA2-7B by a factor of 2.60×2.60$\times$2.60 × while maintaining the average score decrease within 1%, whereas Sheared-LlaMA-1.3B achieves 2.50×2.50$\times$2.50 × with $15.57\%$ performance decreases. As for UNet and MobileNetV3, our method can compress UNet and Pruned-UNet by 7.87×7.87$\times$7.87 × and 7.05×7.05$\times$7.05 × with the performance loss contained in 1%. Particularly, our method succeeds in combination with other model compression methods such as pruning to achieve an even higher compression ratio. In UNet, the total compression ratio is 17.74×17.74$\times$17.74 × with the performance loss $3.41\%$.
No post-hoc retraining. Because discrepancies between the original and decoded parameters are minor, at magnitudes ranging from $10^{-4}$ to $10^{-3}$, the impact on error accumulation through layer-by-layer propagation is acceptable. As a result, the destructive effect of our method is minimal, obviating the need for post-hoc retraining. This advantage is particularly useful in real-world settings, where often 1) training data are inaccessible, and curating data for fine-tuning is costly; 2) no computing resources are supplied for retraining.
Affordable inference time. A network is a hierarchical structure. To expedite the inference time, our approach is to leverage this hierarchy by parallelizing the decoding and inference processes. The key is to complete the decoding of a layer before using this layer to infer. As shown in Table 2, “Original" refers to the time required for a single inference of the original UNet, and “Ours" refers to the time required for one inference by employing parallel decoding and inference processes. It can be seen that the hyper-compression only increases the total inference time moderately.
Short compression time. As Table 6 shows, our method can compress models very fast. This efficiency primarily stems from using matrix operations for most compression processes and the K-D tree to expedite the search. Additionally, by treating the compression tasks between layers as independent operations, we implement a multiprocessing parallel strategy to further decrease the compression time.
Materials and Methods
Motivation
A small human genome can remarkably encode the development of a human brain that has finally billions of neurons and trillions of synaptic connections [23, 22]. This observation suggests the existence of an implicit mapping from genotype to phenotype, capable of expanding a limited number of genetic instructions into a vast array of biological substances. Inspired by this efficient genetic encoding, the hypernet uses a small network (genotype, called hypernet) to generate weights for the target network (phenotype).
Figure 1: We use a parametric function $W_{n}=h(n;\Theta)$ to encode the relationship between locations and weights of the target network, where $W_{n}$ is the $n$-th parameter of the target network. $h$ is implied by the ergodic theory.
We consider the hypernet in the setting of model compression. Furthermore, we generalize the idea of hypernet to a parameterized function (hyperfunction). The biological observation is the existence of a mapping, regardless of whether it is denoted by a network or other forms of functions. With a hyperfunction, instead of directly diminishing weights, one stores a hyperfunction about the model and queries its weights when needed. This novel perspective converts the model compression problem into the parameter representation issue. Mathematically, we use a hyperfunction $W_{n}=h(n;\Theta)$ to encode the relationship between locations and weights of the target network, where $W_{n}$ is the $n$-th parameter of the target network, and $\Theta$ is the hyperparameters of $h$. Along this direction, the important question is how can we design a suitable hyperfunction?
Tools from Ergodic Theory
The above question fits the scope of Diophantine approximations or ergodic theory. A branch of ergodic theory [6] studies whether a low-dimensional system’s trajectory can cover all points in a high-dimensional space. From an engineering perspective, it assesses the feasibility of using low-dimensional curves to approximate high-dimensional points, indicating the potential for model compression. The following theorem illustrates a simple yet notable example from ergodic theory.
Theorem 1.
Suppose $a_{1},\cdots,a_{N}$ are irrationally independent, for any given set $\{W_{n}\}_{n=1}^{N}\subseteq[0,1]$ and $\epsilon>0$, there exists a value $\theta^{*}\in[0,+\infty)$ such that
$|W_{n}-\tau(\theta^{*}a_{n})|<\epsilon,\;n=1,\cdots,N,$ (1)
where $\tau(x)=x-\lfloor x\rfloor$.
The proof of Theorem 1 is a direct consequence of Lemma 19 from [32]. In this theorem, the irrationality independence condition is readily fulfilled; for instance, one can select $a_{k}=1/(\pi+k)$. Theorem 1 provides a natural and elegant way to encode a group of numbers into one number. However, we should only be cautiously optimistic about its application in compression, since simply stitching digits of numbers can also ensemble a group of numbers into one number. What matters is whether we can find a simple $\theta$ that approximates $W_{1},\cdots,W_{K}$ based on (1), such that the memory footprint of the former is smaller than the latter. Our experiments empirically confirm that the parametric representation according to Theorem 1 has compression efficacy.
Core Issues in Algorithmic Development
The entire algorithmic development is anchored in (1). The core objective of the algorithm is, given $\alpha_{1},\cdots,\alpha_{K}$ and an acceptable error $\epsilon$, to identify the smallest $\theta^{*}$ for points in $[0,1]^{K}$ and the highest feasible dimension $K$. First, because a large model usually has billions of parameters, searching $\theta^{*}$ for a group of $K$ parameters has to be fast. Second, an imprecise $\theta^{*}$ could lead to significant performance degradation in the restored networks. Therefore, the algorithm must ensure that $\theta^{*}$ assists in approximating a network’s parameters precisely. Third, when time and precision allow, the algorithm should return the smallest $\theta^{*}$ for the highest compression ratio.
Conclusion and Future Work
In this study, we have proposed hyper-compression, a novel and general-purpose methodology for model compression that leverages hyperfunction to encode the parameters of the target network. We have conducted experiments on various architectures, including LlaMA2-7B, UNet, and MobileNetV3, demonstrating that our approach is both user-friendly and scalable. Nevertheless, it represents merely an initial foray into the whole blueprint. Future directions include 1) theoretically describing the relationship among the error, the length of $\theta$, and the dimension $K$ to reveal the essence of using the ergodic theory to realize compression; 2) beyond classical theory, exploring other possibilities such as implicit neural representation [3] to strike a better balance between compression time, inference time, and compression ratio. We believe that hyper-compression can contribute to Moore’s law of model compression in the near future, i.e., the compression efficiency can be doubled annually, as a solution for the stagnation of hardware Moore’s law.
2 Supplementary Information
I. The Enlarged Gap Between Models’ Size and Hardware Capacity
In recent years, the field of AI has witnessed remarkable advancements, particularly with the development of large language models (LLMs). These large models have demonstrated impressive capabilities in various tasks, which was widely believed to make a big stride towards artificial general intelligence. Large models have rapidly increased in size over the past few years. For instance, GPT-3 consists of 175 billion parameters. Recently, PanGu-$\Sigma$, one of the largest models to date, staggeringly has over $10^{3}$ billion parameters. Such a growth in model size has been driven by the desire to improve the model’s complex language understanding by scaling law. On the other hand, the development of GPU computing capacity such as memory has not kept pace. One main reason is that the fabrication of more advanced GPUs gradually gets closer to the physical ceiling. The disparity between models’ size and GPU memory has created significant challenges in the development of large models.
Figure 2: The rapid growth of large models’ sizes has outpaced the growth of GPU memory, creating huge challenges in serving large models.
Specifically, the disparity between the model size and GPU capacity brings the following problems:
  1. 1.
    Memory Bottleneck: The limited GPU memory poses a bottleneck for serving LLMs. When deploying LLMs for real-world applications, the entire model and its associated parameters need to fit within the GPU memory. However, as LLMs continue to grow in size, it becomes increasingly difficult to accommodate them within the limited memory constraints. This limitation hampers the deployment and wide adoption of LLMs in the edge settings.
  2. 2.
    Ethic Issues: Since large models are hard to serve, large institutes will dominate the development and deployment of large language models, which upsets the whole AI ecosystem:
    Algorithmic Bias: LLMs developed by dominant institutes may be biased by the data they were trained on. This can perpetuate existing societal biases and discrimination, leading to unfair outcomes in decision-making processes.
    Lack of Transparency: Large institutes may not always disclose the full details of their model architectures, training data, or decision-making processes. The lack of transparency can hinder accountability, making it difficult to assess the fairness and reliability of LLMs in real-world applications.
    Ethical Use Cases: The ethical implications of deploying LLMs in various applications, such as content moderation, surveillance, or healthcare, need to be carefully considered. Large institutes’ control over these technologies raises concerns about how they are used and the potential impact on individuals and society.
As the demand for more powerful models and the risk of monopoly by large institutes continue to rise, addressing the disparity between the model size and GPU capacity becomes increasingly important. Particularly, we need to cultivate novel model compression techniques that can realize "Moore’s law of model compression", as a solution for the stagnation of hardware Moore’s law. Once accomplished, Moore’s law of model compression, which proposes that the compression ratios of models should double annually, can constitute a solid infrastructure to support the whole AI ecosystem.
II. Limit of the Compression Ratio of Pruning Convolutional Networks
In the main body, we mention that pruning-based methods are hard to scale. Here, we show that under certain conditions, the compression ratio of pruning a convolutional network is constantly bounded. The network is a randomly generated network with infinite width and depth. We introduce the percolation theory [12] that characterizes the connectivity of a network when links are removed per certain rules.
We study one-dimensional convolutional networks whose topology is shown in Figure 6. We assume any edge of a network is randomly pruned with the same probability. This is also equivalent to the following pruning process: Randomly generate weights for edges of a given network. If the weight of an edge is higher than a threshold, the edge is retained; otherwise, it is removed. In the jargon of percolation theory, an edge is called open or closed if the edge exists or disappears. The percolation theory is a well-known phase transition theory. When the probability of removing edges reaches a threshold, referred to as the percolation threshold, there will be no connected clusters from the input and output, meaning that the network will always yield 00 for any input. The probability threshold directly indicates the maximal proportion of a network’s edges to prune. This section discusses the upper and lower bounds of the percolation threshold for a class of graphs defined by the one-dimensional convolutional network, thereby quantitatively characterizing the compression ratio.
Let us first provide relevant definitions below:
Definition 1 (Graph, Vertex, Edge).
A graph $G=(V,E)$ is an ordered pair of sets, where $V$ is the vertex set, and $E$ is the edge set. In the edge set $E$, we present the edge between $v,w\in V$ as $\langle v,w\rangle$.
The class of graphs we investigate is $G_{r}=(V_{r},E_{r}),~{}r\in\mathbb{Z}_{\geq 2}$, where $V_{r}=\mathbb{Z}^{2},E_{r}=\{\langle(m,n),(m+1,n+i)\rangle\mid m,n\in\mathbb{Z},i\in{0,1,\cdots,r-1}\}$, that is, a convolution with the kernel size $r$ and stride 1. Figure 6 and 6 show the cases of $r=2$ and $r=3$.
Definition 2 (Open and Closed).
Let $p\in[0,1]$. Declare each edge of $G_{r}$ to be open with probability $p$ and closed otherwise, independently of all other edges.
Definition 3 (Open Cluster).
Consider the random subgraph of $G_{r}$ containing the vertex set and the open edges only. The connected components of this graph are called open clusters. We write $C_{r}(v)$ for the open cluster containing the vertex $v$, and we call $C_{r}(v)$ the open cluster at $v$.
Definition 4 (Percolation probability and critical probability).
Let $\theta_{r}(p)$ being the probability that given vertex, $v$, belongs to an infinite open cluster in $G_{r}$, that is $\theta_{r}(p)=P(\mid C_{r}(v)\mid=\infty)$. By the translation invariance of those graphs, we lose no generality by taking this vertex as the origin and denoting $C_{r}(0,0)$ as $C_{r}$. The critical probability is defined to be $p_{c,r}=\sup\{p\in[0,1]\mid\theta_{r}(p)=0\}$.
Earlier, researchers have mainly focused on planar graphs, but we observe that $G_{r}$ is a planar graph only when $r=2$, and is non-planar for $r\geq 3$, parts of $G_{2}$ and $G_{3}$ are shown in Figures 6 and 6. Therefore, many methods, such as using the duality of planar graphs and the star-triangular transition, cannot be applied. In the following, we will first solve the critical probability for the case of $r=2$, and then study the upper and lower bounds of the critical probability for $r\geq 3$. Our method is to degenerate a non-planar graph into a planar graph by removing edges.
Figure 3: $G_{2}$
Figure 4: $G_{3}$
Figure 5: $G_{-1}$
Figure 6: $G_{triangular}$
Proposition 1 (Critical Probability of $G_{2}$).
$p_{c,2}=0.5$.
Proof of Proposition 1.
Let $\mathbb{L}^{2}$ be the square lattice on $\mathbb{Z}^{2}$, whose percolation threshold is well-known to be 0.5 [12]. Consider $\phi:\mathbb{Z}^{2}\longrightarrow\mathbb{Z}^{2}$ defined by $\phi(m,n)=(m-n,n)$. We are going to show $\phi$ is graph isomorphism.
For points $(m,n),(m+1,n)$ connected by the edge $\langle(m,n),(m+1,n+i)\rangle$ in $G_{2}$, we have $\phi(m,n)=(m-n,n),\phi(m+1,n+i)=(m-n+1,n)$ are connected by $\langle(m-n,n),(m-n+1,n)\rangle$ in $\mathbb{L}^{2}$.
For points $(m,n),(m+1,n+1)$ connected by the edge $\langle(m,n),(m+1,n+1)\rangle$ in $G_{2}$, we have $\phi(m,n)=(m-n,n),\phi(m+1,n+1)=(m-n,n+1)$ are connected by $\langle(m-n,n),(m-n,n+1)\rangle$ in $\mathbb{L}^{2}$.
Therefore, $\phi$ is a graph isomorphism from $G_{2}$ to $\mathbb{L}^{2}$, and hence the percolation threshold of $\mathbb{L}^{2}$ is the same as that of $G_{2}$, which is 0.5.
Proposition 2 (Critical Probability of $G_{r},r\geq 3$).
For $r\geq 3$, $\frac{1}{\mu_{r}}\leq p_{c,r}\leq p_{0}$, where $\mu_{r}$ is the connected constant of $G_{r}$ and $p_{0}$ is the solution of $12p+p^{2}-p^{4}=$ 1 between 0 and 1.
Proof of Proposition 2.
The proof of the part $\frac{1}{\mu_{r}}\leq p_{c,r}$ is a more general result that the percolation threshold of any graph is no less than the reciprocal of its connective constant, see [12, 2].
The proof of the part $p_{c,r}\leq p_{0}$ is even more complicated. Define a graph $G_{-1}=(V_{-1},E_{-1})$, where $V_{-1}=\mathbb{Z}^{2},E_{-1}=\{\langle(m,n),(m+1,n+i)\rangle~{}\mid~{}m,n\in\mathbb{Z},~{}i\in\{-1,0,1,\cdots,r-2\}\}$, and Figure 6 shows the $G_{-1}$ corresponding to $G_{3}$. Meanwhile, consider the function $\Phi:\mathbb{Z}^{2}\longrightarrow\mathbb{Z}^{2}$ defined by $\phi(m,n)=(m,n-m)$, which maps the two points $(m,n),(m+1,n+i)$ connected by an edge in $G_{r}$ to $(m,n-m),(m-n,n-m+i-1)$ that are connected by an edge in $G_{-1}$. Thus, $\Phi$ is a graph isomorphism from $G_{r}$ to $G_{-1}$, and they have the same percolation threshold, i.e., $p_{c,r}=p_{c,-1}$.
Next, we modify $G_{-1}$ by removing or connecting vertices and edges.
Remove vertices: $(2m+1,2n),(2m,2n+1),m,n\in\mathbb{Z}$,
Remove edges: $\langle(2m+1,2n),(2m+2,2n+i)\rangle,\langle(2m,2n+1),(2m+1,2n+j),\rangle,i=-1,1,\cdots,r-2,j=0,2,\cdots,r-1$.
Connect pair of edges $\langle(2m,2n),(2m+1,2n)\rangle$ and $\langle(2m+1,2n),(2m+2,2n)\rangle$ into $\langle(2m,2n),(2m+2,2n)\rangle$.
Connect pair of edges $\langle(2m,2n+1),(2m+1,2n+1)\rangle$ and $\langle(2m-1,2n+1),(2m,2n+1)\rangle$ into $\langle(2m-1,2n+1),(2m+1,2n+1)\rangle$.
The resulting graph is a triangular lattice (rotated 45 degrees counterclockwise to the usual form), as shown in Figure 6, denoted as $G_{triangle}$. During the construction of $G_{triangle}$, each connected subgraph in $G_{triangle}$ can uniquely correspond to a connected subgraph in $G_{-1}$, i.e., those untouched edges and vertices are naturally embedded, those connected edges are restored to one vertex and two edges.
To maintain the correspondence with $G_{-1}$ probabilistically, the edges formed by merging in $G_{triangle}$ have probability $p^{2}$, while the remaining edges have probability $p$ of being open. Under the probability, the probability that each connected subgraph in $G_{triangle}$ is open is the same as the probability that its corresponding subgraph in $G_{-1}$ is open. Furthermore, each infinite open cluster in $G_{triangle}$ can uniquely correspond to an infinite open cluster in $G_{-1}$, but not vice versa. Therefore, $\theta_{-1}(p)=P(\mid C_{-1}(v)\mid=\infty)\geq P(\mid C_{triangle}(v)\mid=\infty)=\theta_{triangle}(p)$, and thus $p_{c,-1}\leq p_{0}$.
Fortunately, this inhomogeneous triangular lattice $G_{triangle}$ has been studied extensively [12], and the solution under the provided three-edge probabilities is the solution of $12p+p^{2}-p^{4}=$ 1 between 0 and 1, which is approximate 0.425787.
Remark. From the viewpoint of percolation theory, the compression ratio of pruning a randomly generated convolutional network is $1/p_{0}<\frac{1}{p_{c,r}}<\mu_{r}$, which is a fixed constant given $r$. Our analysis is based on an assumption that ignores the association between two edges. In reality, when an edge with a larger weight is pruned, edges with smaller weights are also pruned. Future work will be considering the association to have a more accurate estimation for the pruning bound.
III. Algorithmic Details
Anchored in the hyperfunction driven by ergodic theory mentioned in the main body, we design a general-purpose model compression method that enjoys the following merits (PNAS): Preferable compression ratio, No post-hoc retraining, Affordable inference time, and Short compression time. In this section, we describe in detail our entire model compression algorithm which is made of four parts, and each part corresponds to one facet of merits.
Figure 7: An illustrative scheme for using ergodic theory to learn a low-dimensional representation of high-dimensional points.
Compression via Ergodicity $\to$ Preferable compression ratio
The core objective of the algorithm is to identify the smallest $\theta^{*}$ for points in hypercubes $[0,1]^{K}$ under the condition that $\epsilon$ is sufficiently small and $K$ is as high as possible. To this end, we consider the distribution of weights of the original network by binning weights into different groups based on magnitudes of weight values.
  • First, we flatten all parameters into a one-dimensional vector $\mathbf{W}=[W_{1},W_{2},\cdots,W_{N}]$, where $N$ is the total number of weights, and then split it into groups $\left[\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}\right]$, where $W^{(g)}=[W_{K(g-1)+1},W_{K(g-1)+2},\cdots,W_{K(g-1)+K}],g=1,2,\cdots,G$ is of $\mathbb{R}^{1\times K}$, and $G$ is the number of groups. For any $\mathbf{W}^{(g)}$, there exists an integer $\theta^{*}_{g}$ such that $|\mathbf{W}^{(g)}-\tau(\theta^{*}_{g}\cdot\mathbf{a})|<\epsilon,g=1,\cdots,G$ and $\mathbf{a}=[a_{1},a_{2},\cdots,a_{K}]$. Finally, the vector $\left[\theta^{*}_{1},\theta^{*}_{2},\cdots,\theta^{*}_{G}\right]$ represents a compression to $\mathbf{W}$. Take $K=2$ and $\mathbf{W}=[W_{1},W_{2},W_{3},W_{4}]=[0.1,0.2,0.4,0.5]$ as an example, we split $\mathbf{W}$ as $[\mathbf{W}^{(1)},\mathbf{W}^{(2)}]$, where $\mathbf{W}^{(1)}=[W_{1},W_{2}]=[0.1,0.2]$ and $\mathbf{W}^{(2)}=[W_{3},W_{4}]=[0.4,0.5]$, and derive $\theta^{*}_{1}$ and $\theta^{*}_{2}$ for $\mathbf{W}^{(1)}$ and $\mathbf{W}^{(2)}$, respectively. As Figure 7 shows, we can find a $\theta$ such that $(\tau(\theta/(\pi+1)),\tau(\theta/(\pi+2)))$ can approximate any point in 2D space.
[…]
  • In our algorithm, we set $K=2$ and replace the unit square $[0,1]^{2}$ with a more flexible box $[a,b]\times[c,d]$. Specifically, given $\mathbf{W}$, $[a,b]\times[c,d]$ is defined as a square with side length $l$, centered at $(\bar{x_{i}},\bar{y_{i}})$, where $(\bar{x_{i}},\bar{y_{i}})$ is the centroid of two-dimensional points $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$. $l$ is a hyperparameter typically set to 0.01 for optimal results. Consequently, the function $\tau(x):=x-\lfloor x\rfloor$ in the ergodic theorem is generalized as $\tau([x,y]):=f([x,y])+([\bar{x_{i}},\bar{y_{i}}]-[\frac{l}{2},\frac{l}{2}])$, where $f(x)=x\bmod l$.
  • If we construct a square with its center at the centroid $(\bar{x_{i}},\bar{y_{i}})$ and the side length $l$, it is highly probable that many points in $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$ will fall outside this square. Therefore, we need to first scale these points into the square using a scaling factor $s$ to compress, and then scale the substituted points $\tau(\theta^{*}_{g}\cdot\mathbf{a})$ back during decompression. While the error rate remains intact, this inversion will amplify the error $\epsilon$, where $\epsilon=|W^{(g)}-\tau(\theta^{*}_{g}\cdot\mathbf{a})|$. Assuming $\mathbf{W}^{(F)}$ is the point farthest from the centroid $(\bar{x_{i}},\bar{y_{i}})$ in $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$, we define the farthest distance $l_{f}$ as $|\mathbf{W}^{(F)}-[\bar{x_{i}},\bar{y_{i}}]|$, and then the scaling factor $s=\frac{l}{2\cdot l_{f}}$. The specific formula derivation is as follows:
    $\mathbf{W}^{(in)}$ $=\mathbf{W}^{(in)}-C$ (2)
    $=C\mathbf{W}^{(out)}\cdot s$
    $=C\mathbf{W}^{(out)}\cdot\frac{l}{2\cdot l_{f}},$
    where $C$ is the centroid node, $\mathbf{W}^{(out)}$ is a node ouside the squre, and $\mathbf{W}^{(in)}$ is the scaled node of $\mathbf{W}^{(out)}$.
    We denote $\tau(\theta^{*}_{i}\cdot\mathbf{a})$ as $\mathbf{W}^{(in)*}$, so that $\epsilon=\mathbf{W}^{(in)*}-\mathbf{W}^{(in)}$. The detailed scale-back process is shown as follows:
    $\mathbf{W}^{(out)*}$ $=\frac{C\mathbf{W}^{(in)*}}{s}+C$ (3)
    $=\frac{\mathbf{W}^{(in)}+\epsilon}{s}+C.$
    Therefore, we can derive the error:
    $error$ $=|\mathbf{W}^{(out)}-\mathbf{W}^{(out)*}|=|C\mathbf{W}^{(out)}-C\mathbf{W}^{(out)*}|$ (4)
    $=|C\mathbf{W}^{(out)}-\frac{C\mathbf{W}^{(in)*}}{s}|=|\frac{\mathbf{W}^{(in)*}-\mathbf{W}^{(in)}}{s}|$
    $=|\frac{\epsilon}{s}|=|\frac{2\cdot\epsilon\cdot l_{f}}{l}|$
    From (4), we can observe that the farthest distance $l_{f}$ affects the error of estimating $\mathbf{W}^{(out)}$ after scaling back.
    Figure 8: Given $K=2$, the points outside the center square are divided into three categories: blue points belong to the first category, pink points belong to the second category, and green points belong to the third category.
    Therefore, as shown in Fig. 8 in our algorithm, we classify the points outside the square into $M$ categories based on their distances from the centroid $C$, where $M$ is a hyperparameter, so that the points in different categories can utilize different farthest distance $l_{fm}$ to define different scale factor $s_{m}$. This is better than using the global farthest distance $l_{f}$ to control the compression error. This method is highly effective in practice, as a significant portion of the parameter nodes $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$ exhibit a pattern that most points are close to the centroid.
    Specifically, assuming that the points outside the square are divided into $M$ categories and $W^{(m)}$ is a point under the $m$-th category, the scaling factor $s_{m}$ for this point is defined as
    $s_{m}=\frac{l/2}{l/2+(l_{f}/K)\cdot m}.$ (5)
    The final compression $\theta_{m}^{*}$ is given by the following formula:
    $\theta_{m}^{*}=m\cdot U+\lambda_{m}^{*},$ (6)
    where $\lambda_{m}^{*}$ is the compression of the scaled $W^{(m)}$ in the center square.
Adjusting Irrational Direction $\to$ No fine-tuning
Our method does not require fine-tuning of the model, and can directly use the trained model for compression. This greatly simplifies the model compression process and avoids the time and computing resource consumption caused by retraining the model. One reason accounting for why our algorithm requires no retraining is mainly because our algorithm can guarantee the error rate is low by optimizing the direction of $\mathbf{a}$, the side length $l$, and $U$. The error rate of points in different classes is always no more than $(\frac{l}{U})/l=\frac{1}{U}$.
Specifically, in the ergodic theorem, $\mathbf{a}=[a_{1},a_{2}]=\mathbf{d}\cdot I$ is a two-dimensional directional vector, where $\mathbf{d}$ and $I$ determine the direction vector and step size for the movement of curves defined by the parametric equation, respectively. Since $\theta$ is an integer no more than $U$, at most we have $U+1$ points in two-dimensional space $(\tau(a_{1}\theta),\tau(a_{2}\theta)),\theta=0,1,\cdots,U$. In our experiment, we define an optimal vector $\mathbf{a}$ based on $U$, such that the distribution of $U+1$ points is more uniform, thereby minimizing the maximum error as much as possible. Specifically,
$\mathbf{a}$ $=\mathbf{d}\cdot I$ (7)
$=[\frac{l}{U},l]/\|\frac{l}{U},l\|\cdot I$
$=[\frac{l}{U},l]/\|\frac{l}{U},l\|\cdot\frac{l}{\sin{\alpha}\cdot\lfloor\sqrt{U}\rfloor},$
where $\tan{\alpha}=\frac{l}{l/U}=U$. As shown in Fig. 9, it demonstrates that using (7) results in a more uniform distribution of the sample nodes compared to defining $\mathbf{a}$ as follows:
$\mathbf{a}=[1/(\pi+1),1/(\pi+2)].$ (8)
Figure 9: Different $\mathbf{a}$ will result in different distribution of points. Upper three subfigures use (7), while lower three figures utilize (8).
Conducting Inference and Decompression Simultaneously $\to$ Affordable inference time
At first glance, one may think that model compression based on hyperfunctions suffers from a slow inference time, as this technique needs to restore parameters before inference, which adds another level of computation. Here, we propose a scheme that leverages the intrinsic hierarchical structure of a network to greatly reduce the inference time. Our scheme parallelizes parameter decompression and inference. As shown in Figure 10, while the parameters of later layers (except the first layer) are restored, the inference operation in earlier layers is also carried out simultaneously. As long as we can recover the parameters of the current layer before the inference arrives at this layer, there is no waste of time. Thus, theoretically, the inference time of using our algorithm only increases quite moderately. The increment is the time used to restore the parameters of the first layer.
Figure 10: The rapid growth of LLM’s size has outpaced the growth of GPU memory, creating challenges in serving these increasingly massive models.
Based on the above solution, we need to shorten the decompression time. First, the entire decompression process is implemented using matrix operations, which can significantly speed up the computation. Second, the process of reading files from storage and preprocessing them is relatively time-consuming, and repeats each time the function is called. Therefore, we optimize this process, performing file reading and preprocessing operations only when the function is called for the first time, and storing the processing results in the cache. When the function is called again, the preprocessed intermediate results are directly captured from the cache, further accelerating the inference.
KD Tree + Parallelization $\to$ Short Compression Time
With the advent of large models, compression time becomes an important facet of evaluating a compression algorithm. Here, we propose a suite of techniques to enable the proposed hyper-compression method to fast complete the model compression process with off-the-shelf tools, for both CPU and GPU inference.
  • As mentioned earlier, when encoding the parameter vector $\mathbf{W}$, we first convert it into many points $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$. Then, given $\mathbf{W}^{(g)}$, we search $u$ from $\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})$ to approximate $\mathbf{W}^{(g)}$. Next, instead of repeating searching for every point, we store $\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})$ in the format of k-d trees [19]. Thus, we can fast determine $\tau(u\cdot\mathbf{a})\in\{\tau(1\cdot\mathbf{a}),\cdots,\tau(U\cdot\mathbf{a})\}$ that is closet to $\mathbf{W}^{(g)}$.
  • We extensively use matrix operations. Given a series of two-dimensional point coordinates $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}$, we first calculate the scaling factor list $F$:
    $F$ $=[s_{1},s_{2},\cdots,s_{G}]^{\top},$ (9)
    $s_{g}$ $=\frac{l/2}{l/2+(l_{f}/M)\cdot m_{g}},$
    where $s_{g}\in F$ represents the scale factor of $\mathbf{W}^{(g)}$ and $m_{g}$ means the category of node $\mathbf{W}^{(g)}$. The points located within the center square are classified as category $m_{g}$.
    Let $\mathbf{m}=\left[m_{1},y_{2},\cdots,m_{G}\right]^{\top}$, $\mathbf{W}=\left[\mathbf{W}^{(1)},\mathbf{W}^{(2)},\cdots,\mathbf{W}^{(G)}\right]^{\top}$. Then, we can calcute the final compression $\theta^{*}$ as follows:
    $C\mathbf{W}=\mathbf{W}^{\top}-[C,\cdots,C]^{\top},$ (10)
    $O=C\mathbf{W}\cdot F+[C,\cdots,C]^{\top}$
    $\lambda^{*}=KD(O^{\top},[\tau(1\cdot a),\cdots,\tau(U\cdot\mathbf{a})])$
    $\theta^{*\top}=\lambda^{*\top}+U\cdot\mathbf{m}$
  • We also multiprocess the compression by simultaneously compressing different layers of a model to maximize the utilization of computational resources.
VI. Experimental Details
In this section, we introduce details of experimental configurations. All experiments are conducted on a single NVIDIA RTX4090 (24GB) GPU using the PyTorch framework.
  • Details of experiments related to UNet: We perform parameter compression on both UNet and the pruned UNet, and test the performance loss of the pruned models on the Carvana [1] dataset.
    1. 1.
      UNet: U-Net [20] is a deep convolutional neural network architecture that has proven highly effective for image segmentation tasks, especially in biomedical imaging. The network adopts an encoder-decoder structure. The encoder progressively downsamples feature maps to capture context, while the decoder upsamples these feature maps and concatenates them with corresponding feature maps from the encoder via skip connections to precisely localize objects.
    2. 2.
      Pruned UNet: Specifically, the pruned UNet refers to the UNet model whose parameters are removed using the Taylor-Rank Pruning [18] with the FLOPs regularization [13] (strength=0.001). This is essentially a channel pruning method. The bar chart below illustrates the percentage of channels pruned in each layer of the pruned UNet relative to the original UNet.
      123456789101112131415161718192021220020202020404040406060606080808080100100100100LayerRelative % of Channels Removed
    3. 3.
[…]
    4. 4.
      Evaluation: The Dice coefficient [7], also known as the Dice similarity index, is a statistical measure used to gauge the similarity between two sets. It is particularly popular in the field of image segmentation to assess the accuracy of models by comparing the overlap between the predicted segmentation and the ground truth. Mathematically, the Dice coefficient is defined as
      $\mathtt{Dice}=\frac{2|A\cap B|}{|A|+|B|},$ (11)
      where $A$ represents the predicted segmentation, and $B$ is the ground truth. The value of the Dice coefficient ranges from 00 to $1$, where $1$ indicates the perfect agreement between two sets, and 00 indicates no overlap at all.
  • Details of experiments related to MobileNetV3: We apply our compression method to both MobileNetV3 and the pruned MobileNetV3 optimized by automated gradual pruning algorithm [33], and evaluate the top-1 accuracy on the CIFAR-10 dataset [16].
    1. 1.
      MobileNetV3: MobileNetV3 is a state-of-the-art convolutional neural network architecture designed specifically for mobile devices. Building upon the success of its predecessors, MobileNetV3 introduces novel design principles and optimization techniques to achieve a more compelling balance between accuracy and efficiency.
    2. 2.
      Pruned MobileNetV3: We employ the Automated Gradual Pruner (AGP) [33] technique to prune MobileNetV3, ultimately reducing the model size by approximately 2.68 times. The central idea of Automated Gradual Pruner (AGP) is to prune the model’s weights gradually over time, allowing the network to adapt to the loss of parameters through continuous retraining. This technique is particularly effective in minimizing the performance degradation associated with aggressive pruning strategies.
    3. 3.
      Dataset: The CIFAR-10 dataset [16] is a well-known benchmark in the field of machine learning and computer vision, consisting of $60,000$ $32\times 32$ colored images belonging to $10$ classes. The dataset is split into $50,000$ training images and $10,000$ testing images.
    4. 4.
      Evaluation: The top-1 accuracy is an evaluation metric designed particularly for classification tasks, which primarily tests whether a model can make a correct single prediction for each input sample in a given task. Mathematically, let $\mathbf{x}_{i}$ represent an input instance, and let $\hat{y}_{i}=\arg\max_{j}\;p(y_{j}~{}|~{}\mathbf{x}_{i})$ denote the predicted label, where $p(y_{j}~{}|~{}\mathbf{x}_{i})$ is the predicted probability of the $j$-th class for the input $\mathbf{x}_{i}$. The true label is denoted by $y_{i}$. The top-1 accuracy is defined as the ratio of correctly predicted instances to the total number of instances, and can be expressed as the following formula:
[…]
      where $N$ is the total number of instances, and $\mathbb{I}(\cdot)$ is the indicator function, which returns $1$ if its argument is true and 00 otherwise.
  • Details of experiments related to LlaMA2-7B: We apply our compression method to the LLaMA2-7B [24] and three other performant pruned variants of the LLaMA2-7B: Sheared-LLaMA-1.3B [28], TinyLLaMA [31], and LiteLLaMA. We then examine the performance drop on these models using various metrics, including SciQ [26], WinoGrande [21], ARC-E [5], 25-shot ARC-C [5], 10-shot Hellaswag [30], 32-shot BoolQ [4], 32-shot NQ [17], 5-shot MMLU [14], and Perplexity [27] to evaluate the effects of compression by using the package lm-evaluation-harness [11].
    1. 1.
      LlaMA2-7B [24]: It is a state-of-the-art language model that has 7-billion parameters and was made publicly available by Meta AI. Employing the Transformer architecture, the model has been pre-trained on a massive text corpus, enabling it to exhibit superior performance on a multitude of natural language processing tasks, including text generation, question answering, machine translation, and so on.
    2. 2.
      Other comparative models:
      • Sheared-LlaMA-1.3B [28]: It is a variant of LlaMA2-7B that uses a novel pruning algorithm: targeted structured pruning which can reduce a source model to a specified target architecture defined by the configurations of the existing pre-trained models. Additionally, this algorithm also introduces a dynamic batch loading strategy that proportionally loads training data from each domain based on its rate of loss reduction, thereby optimizing data utilization and facilitating overall performance improvements.
      • TinyLlaMA [31]: TinyLLaMA is a compact and efficient variant of the LLaMA2-7B model. It retains the same architecture and tokenizer as LLaMA2-7B but scales down to a Transformer decoder-only model with 1.1 billion parameters. This model is trained on up to 3 trillion tokens, with a specific focus on the behavior of smaller models when exposed to a substantially larger quantity of training tokens than those predicted by scaling laws.
      • LiteLlaMA: LiteLlaMA is another variant of LlaMA2-7B with 460M parameters which is trained over 1T tokens on data from the RedPajama dataset [10].
    3. 3.
      Evaluation: In this study, most evaluation metrics for large language models (LLMs) are tested using the open-source package lm-evaluation-harness [11] which is a versatile and open-source Python package for benchmarking the ability of large language models (LLMs) across various tasks and datasets. The "($i$)" after the name of an evaluation metric means "$i$-shot".
      • SciQ [26]: SciQ, a dataset of 13.7k multiple choice science exam questions, is a benchmark dataset specifically designed to evaluate a language model’s ability to understand and reason about scientific questions. By presenting the model with carefully crafted questions that require a deep understanding of scientific concepts and principles, SciQ assesses a model’s capacity to perform complex reasoning tasks.
      • WinoGrande [21]: Winogrande, a collection of 44k WSC-inspired problems, is a benchmark dataset specifically designed to evaluate a language model’s ability to understand and reason about common sense knowledge. By presenting the model with carefully constructed sentences requiring the selection of the correct noun to complete a given context, Winogrande assesses a model’s capacity to grasp nuanced semantic and pragmatic relationships.
      • ARC-E & ARC-C (25) [5]: ARC (AI2 Reasoning Challenge), including totally 7,787 natural, grade-school science questions, is a question set, text corpus, and baselines assembled to encourage AI research in advanced question answering, which is partitioned into a challenge set (ARC-C) and an easy set (ARC-E).
      • HellaSwag (10) [30]: HellaSwag, which stands for "Harder Endings, Longer contexts, and Low-shot Activities for Situations With Adversarial Generations," is designed to test commonsense natural language inference (NLI) in the context of physical situations.
      • BoolQ (32) [4]: BoolQ is a reading comprehension dataset of naturally occurring yes/no questions which are challenging and require a wide range of inference abilities to solve.
      • NQ (32) [17]: Natural Questions (NQ), including 307,372 training examples and 7,830 examples for development, contains real user questions issued to Google search, and answers found from Wikipedia by annotators, designed for the training and evaluation of automatic question answering systems.
      • MMLU (5) [14]: Massive Multitask Language Understanding (MMLU) benchmark is designed to evaluate a language model’s ability to generalize across a wide range of tasks and domains. By assessing performance on a diverse set of 57 tasks, including elementary mathematics, US history, computer science, law and more, MMLU provides a comprehensive evaluation of a model’s extensive world knowledge and problem solving ability.
      • Perplexity [27]: Perplexity (PPL) is used to measure the quality of a probability distribution or probabilistic model in predicting samples. A model with lower perplexity indicates better performance in predicting samples.
    4. 4.
      More detailed information on all the models used for comparative testing:
      Model Parameter Precision Number of Parameters File Size
      UNet [20] FP32 13,395,329 52.4MB
      Pruned-UNet FP32 5,326,021 20.3MB
      MobileNetV3 [15] FP32 1,528,106 5.95MB
      Pruned-MobileNetV3 FP32 556,442 2.22MB
      LlaMA2-7B [24] FP16 6,738,415,616 12.50GB
      Sheared-LlaMA-1.3B [28] FP32 1,345,423,360 5.01GB
      TinyLlaMA [31] FP32 1,100,048,384 4.09GB
      LiteLlaMA  [^1] For more details, please visit the Hugging Face website of LiteLlaMA: https://huggingface.co/ahxt/LiteLlama-460M-1T. FP16 461,685,760 0.86GB
III. Supplementary Experimental Results
In this section, we present supplementary experimental results that are not put into the main body due to the limit of pages.
2.1 Compression Performance
Table 4: Comparison of Our Methods with Other Compression Variant Models of LlaMA2-7B on Individual Evaluation Dataset
Model File Size (GB) SciQ (%) WinoGrande (%) ARC-E (%) ARC-C (25) (%)
LlaMA2-7B [24] 12.50 94.00 68.98 76.30 52.39
Sheared-LlaMA-1.3B [28] 5.01 (2.50$\times$) 87.30 58.09 60.98 34.04
TinyLlaMA [31] 4.09 (3.06$\times$) 89.30 59.43 61.66 37.12
LiteLlaMA  [^1] For more details, please visit the Hugging Face website of LiteLlaMA: https://huggingface.co/ahxt/LiteLlama-460M-1T. 0.86 (14.53$\times$) 75.10 52.64 47.85 24.32
LlaMA2-7B + HF 4.80 (2.60$\times$) 93.70 69.77 75.59 53.24
Sheared-LlaMA-1.3B + HF 0.98 (12.76$\times$) 87.90 58.88 60.48 32.85
TinyLlaMA + HF 0.78 (16.03 $\times$) 89.50 58.96 61.11 36.43
LiteLlaMA + HF 0.39 (32.05$\times$) 73.00 54.22 44.78 23.89
Model HellaSwag (10) (%) BoolQ (32) (%) NQ (32) (%) MMLU (5) (%) Average (%)
LlaMA2-7B [24] 78.94 81.90 28.67 45.86 65.88
Sheared-LlaMA-1.3B [28] 61.02 65.54 9.89 25.59 50.31
TinyLlaMA [31] 62.48 62.91 12.52 26.79 51.53
LiteLlaMA 38.41 57.09 1.77 26.11 40.41
LlaMA2-7B + HF 77.17 80.92 25.65 43.04 64.89
Sheared-LlaMA-1.3B + HF 60.56 63.76 8.98 24.68 49.76
TinyLlaMA + HF 61.92 58.41 11.58 27.28 50.65
LiteLlaMA + HF 37.56 56.57 1.16 26.61 39.72
Table 5: The comparison of Perplexity (PPL) on the dataset wikitext-2-raw-v1
Model Compression Ratio PPL
LlaMA2-7B - 5.47
LlaMA2-7B + W8A8 SmoothQuant [29] (int8, $\alpha$=0.85) 2.00$\times$ 5.52
LlaMA2-7B + Asymmetric Quant (int8) 2.00$\times$ 5.65
LlaMA2-7B + GPTQ [9] (int4) 3.46$\times$ 5.69
LlaMA2-7B + Asymmetric Quant (int4) 4.00$\times$ 26160.34
LlaMA2-7B + HF 2.60$\times$ 5.82
Sheared-LlaMA-1.3B 2.50$\times$ 8.13
Sheared-LlaMA-1.3B + HF 12.76$\times$ 8.37
TinyLlaMA 3.06$\times$ 7.71
TinyLlaMA + HF 16.03$\times$ 7.95
LiteLlaMA 14.53$\times$ 31.82
LiteLlaMA + HF 32.05$\times$ 37.85
Figure 11: The exemplary outputs from LlaMA2-7B, Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA, and their compressed model by using hyper-compression. These results demonstrate that our compression method effectively preserves the models’ ability to produce meaningful and grammatically correct text.
Here, we show the compression performance of LLaMA2-7B and seven different variants based on LlaMA2-7B. First, in Table 4, we test eight important downstream tasks. By directly compressing LLaMA2-7B using our method (LLaMA2-7B + HF), a compression ratio of 2.6$\times$ is achieved with only a 0.99% decrease in the average score. When our method is combined with other LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, and LiteLlaMA) respectively, higher compression ratios can be achieved, with the decrease in the average score remaining within 1%, compared to the pruned ones. Moreover, Perplexity (PPL) is an important metric for evaluating the overall performance of large language models (LLMs). We conduct tests on the publicly available dataset wikitext-2-raw-v1222For more details, please visit the website of dataset wikitext-2-raw-v1: https://huggingface.co/datasets/Salesforce/wikitext.. As shown in Table 5, the PPL of LLaMA2-7B+HF increases by only 0.35 compared to the original LLaMA2-7B model, whereas the PPL increases caused by the other three LLaMA2-7B variants (Sheared-LlaMA-1.3B, TinyLlaMA, LiteLlaMA) are 2.66, 2.24, and 26.35, respectively. This demonstrates that our method not only enhances the compression ratio but also effectively maintains the model’s overall performance. We also compare the perplexity with four quantization methods: W8A8 SmoothQuant (int8, $\alpha$=0.85), Asymmetric Quant (int8), GPTQ (int4), and Asymmetric Quant (int4). Among them, both W8A8 SmoothQuant (int8, $\alpha$=0.85) and GPTQ (int4) require calibration sets for calibration during the quantization process, whereas our method does not involve calibration. Figure 11 presents two examples comparing the text outputs generated by eight different LLM models for a given prompt. The experimental data demonstrate that our method achieves a preferable compression ratio while having minimal impact on the model’s performance across various downstream tasks. Moreover, it can also integrate effectively with other compression methods such as pruning.
2.2 Compression Time
As shown in Table 6, we record how much using our method to compress eight models. The time to compress smaller models like UNet and MobileNetV3 is only 35 seconds, with Pruned-MobileNetV3 requiring only 8.83 seconds. The compression time for large language models is longer. For example, compressing LLaMA2-7B takes 53 minutes, while LiteLLaMA is compressed in just 6 minutes. The smaller the model’s size, the faster the compression is.
Table 6: The compression time of our method on UNet, MobileNetV3
and LlaMA2-7B
Model Compression Time (s)
UNet + HF [20] 30.00
Pruned-UNet + HF 34.60
MobileNetV3 + HF 11.63
Pruned-MobileNetV3 + HF 8.83
Model Compression Time (min)
LlaMA2-7B + HF 53
Sheared-LlaMA-1.3B + HF 7
TinyLlaMA + HF 11
LiteLlaMA + HF 6
As shown in Figure 7, we conduct ablation experiments based on UNet and MobileNetV3 on the aforementioned two acceleration techniques to demonstrate their effectiveness in enhancing computational speed.
Table 7: The results of ablation study on acceleration techniques
Model K-D Tree Matrix Operations Compression Time (min)
UNet [20] 231.6
176.9
54.8
0.5
MobileNetV3 [15] 52.29
3.13
45.50
0.19
Furthermore, we show that the concurrent application of k-d trees and matrix operations leads to a substantial enhancement in computational efficiency.
2.3 Inference Time
We test the inference time and the execution time of our method under different batch sizes on NVIDIA GeForce RTX 4060 and NVIDIA A40 graphics cards, respectively.
Table 8: Timing results for different batch sizes on two devices using two methods with a UNet network.
Batch Size 4060 A40
Inference(s) Our Method(s) Inference(s) Our Method(s)
1 4.35 4.43 1.50 2.06
2 3.95 4.49 1.44 2.05
4 3.63 4.77 1.45 2.22
8 3.67 7.94 1.44 2.46
When performing the time test, the model is first initialized and loaded onto the GPU. Next, the start time is recorded, and the model evaluation function is executed 10 times in a loop. After each loop, the CUDA cache is manually cleared and Python’s garbage collection mechanism is triggered to ensure that the memory resource is fully released. After the loop is finished, the end time is recorded, and the average time for each evaluation is calculated.
The Carvana dataset is used in the model evaluation function. This dataset is used for car image segmentation tasks and is widely used in image segmentation benchmarks. Each time the model is evaluated, only 1 batch of data is used for a test. We notice that when batch size increases, the inference time of our compressed model gets closer to that of the original model, which aligns with our expectations. This is because when batch size goes up, the inference time across layers is slow, which allows weights in later layers to be decompressed in time.
V. Parameter Sensitivity
In this section, we use UNet as an example to conduct a parameter sensitivity test on three aforementioned hyperparameters: $M$, $U$, and $l$ in our compression algorithm, where $M$ means the maximum value of categories that can be set for all layers of the model, $U$ means the list of the number of sample nodes in the center square, and $l$ means the length of the side of the center square. As shown in Table 9, it can be observed that different values of $M$, $U$, and $l$ only moderately impact the model’s performance and compression ratio, which means our algorithm is robust to hyperparameters.
Table 9: Parameter sensitivity results for different hyper-parameters on UNet.
[…]
Max Class $U$ $l$
UNet - - - 99.86 -
1 225 0.02 99.81 7.09$\times$
2 225 0.02 97.07 7.08$\times$
3 225 0.02 90.11 7.08$\times$
1 225 0.1 99.21 7.92$\times$
2 225 0.1 96.46 7.88$\times$
3 225 0.1 99.71 7.88$\times$
1 361 0.02 99.87 6.38$\times$
2 361 0.02 97.90 6.38$\times$
3 361 0.02 99.39 6.38$\times$
1 361 0.1 99.28 7.73$\times$
2 361 0.1 99.10 7.71$\times$
3 361 0.1 99.65 7.68$\times$
1 225 / 361 0.02 99.89 7.08$\times$
2 225 / 361 0.02 98.12 7.08$\times$
3 225 / 361 0.02 99.58 7.08$\times$
1 225 / 361 0.1 99.08 7.73$\times$
2 225 / 361 0.1 98.85 7.71$\times$
3 225 / 361 0.1 99.51 7.68$\times$